OpenStudy (anonymous):

A water bed that is 1.5 m wide and 2.5 m long weighs 1055 N. Assuming the entire lower surface of the bed is in contact with the floor, what is the pressure the bed exerts on the floor? a. 250 Pa c. 270 Pa b. 260 Pa d. 280 Pa

OpenStudy (anonymous):

Need Help ASAP

OpenStudy (matt101):

Pressure equals force per area: P=F/A. What's the force here, and over what area is it acting? Find that information and you can find your pressure!

OpenStudy (shamim):

Given F=1055N A=1.5*2.5m^2 p=?